# The number of ends of big mapping class groups

### v1

@@ 2.6. Avenue surfaces and translatable surfaces
Definition 2.7 (Avenue surface).
An avenue surface is a connected orientable surface $\Sigma$ which does not contain any non-displaceable subsurface of finite type, whose end space is tame, and whose mapping class group $\operatorname{Map}(\Sigma)$ is CB-generated but not CB.
The condition of tameness only requires that certain ends of $\Sigma$ admit a stable neighborhood. Since we will eventually only consider stable surfaces, where every end admits a stable neighborhood, we refer the interested reader to [MannRafi] for a precise definition of tame and omit writing it here.
Definition 2.8 (Translatable surface).
Given a surface $\Sigma$, an end $e$ of $\Sigma$, and a sequence $\alpha_{n}$ of simple closed curves on $\Sigma$, we write $\lim_{n\to\infty}\alpha_{n}=e$ if for each neighborhood $V$ of $e$, all but finitely many $\alpha_{n}$ are contained in $V$ (after isotopy). A homeomorphism $h\in\operatorname{Homeo}(\Sigma)$ is a translation if there are two distinct ends $e_{+}$ and $e_{-}$ of $\Sigma$ such that for any simple closed curve $\alpha$ on $\Sigma$, $\lim_{n\to\infty}h^{n}(\alpha)=e_{+}$ and $\lim_{n\to\infty}h^{-n}(\alpha)=e_{-}$. A surface $\Sigma$ is translatable if it admits a translation.
By [Schaffer-Cohen, Theorem 5.3], if a surface $\Sigma$ has tame end space and $\operatorname{Map}(\Sigma)$ is CB generated and not CB, then $\Sigma$ is an avenue surface if and only if it is translatable. In particular, every stable avenue surface is translatable.
@@
Next we recall a characterization of translatable surfaces from [Schaffer-Cohen]. Let $S$ be any surface with exactly two boundary components, and denote by $S^{\natural\mathbb{Z}}$ the surface obtained by arranging countably many copies of $S$ like $\mathbb{Z}$ and then gluing consecutive boundary components together in the natural way. By construction, this surface is translatable. Conversely, every translatable surface admits such a decomposition.
@@ 2.7. The translatable curve graph
We now recall the definition of the translatable curve graph from [Schaffer-Cohen]. Let $\Sigma=S^{\natural\mathbb{Z}}$ be a translatable surface and assume that $\Sigma$ is tame. By [Schaffer-Cohen, Lemma 4.4] and [MannRafi, Proposition 4.7], the end space ${\mathrm{End}}(S)$ has a finite, positive number of equivalence classes of maximal ends, and each such equivalence class intersects ${\mathrm{End}}(S)$ at either a finite set or a Cantor set. So we may fix, once and for all, a set $\mathcal{R}=\left\{{f_{1},\dots,f_{N},c_{1},\dots,c_{M}}\right\}$ of representatives of the equivalence classes of maximal ends of $S$ such that each $E(f_{i})\cap{\mathrm{End}}(S)$ is finite and each $E(c_{i})\cap{\mathrm{End}}(S)$ is a Cantor set. Note that in the whole end space ${\mathrm{End}}(\Sigma)$, since $\Sigma=S^{\natural\mathbb{Z}}$, each $E(f_{i})$ is homeomorphic to $\mathbb{Z}$ and each $E(c_{i})$ is homeomorphic to the Cantor set with two points removed. We say an end of $\Sigma$ is of discrete type if it is equivalent to some $f_{i}$. Observe that $\Sigma$ has at least one end of discrete type if and only if $N>0$.
@@
Pick mutually disjoint stable neighborhoods $V_{f_{i}}$ and $V_{c_{i}}$ of the elements of $\left\{{f_{1},\dots,f_{N},c_{1},\dots,c_{M}}\right\}$. By the definition of stable, each $V_{f_{i}}$ contains exactly one discrete-type end. For each $i=1,\dots,N$, let $T_{i}$ be a subsurface of $S$ with two boundary circles and end space homeomorphic to $V_{f_{i}}\sqcup\bigsqcup_{j=1}^{M}V_{c_{j}}$. If $f_{i}$ or any of the $c_{j}$ is accumulated by genus, then $T_{i}$ necessarily has infinite genus; otherwise, we require $T_{i}$ to have genus 0. If $S$ has finite positive genus, then let $T_{N+1}$ be a subsurface with two boundary circles, genus 1, and end space homeomorphic to $\bigsqcup_{j=1}^{M}V_{c_{j}}$. Put $\mathcal{S}=\left\{{T_{1},\dots,T_{N},T_{N+1}}\right\}$ (including $T_{N+1}$ only if it is defined, of course). The translatable curve graph $\mathcal{T}\mathcal{C}(\Sigma)$ of $\Sigma$ is the graph whose vertices are the isotopy classes of separating curves on $\Sigma$, with an edge between two curves $\alpha$ and $\beta$ if they have disjoint representatives and $\alpha,\beta]$ is homeomorphic to a subsurface in $\mathcal{S}$. By [Schaffer-Cohen, Lemma 4.7], $\mathcal{T}\mathcal{C}(\Sigma)$ is connected, and we denote its path metric by $d$. Observe that $\operatorname{Map}(\Sigma)$ naturally acts by isometries on $\mathcal{T}\mathcal{C}(\Sigma)$.
@@
There is an edge case: if $N=0$ and the genus of $S$ is either 0 or infinite, then the above construction gives $\mathcal{S}=\emptyset$. This is not desirable, so in this case we put $\mathcal{S}=\{S\}$. In some sense this choice is consistent with the construction above because $S$ is indeed a subsurface with two boundary circles and end space homeomorphic to $\bigsqcup V_{c_{i}}$. More importantly, in this case, $\mathcal{T}\mathcal{C}(\Sigma)$ has diameter 2 and $\operatorname{Map}(\Sigma)$ is CB, which means that both are quasi-isometric to a point (see [Schaffer-Cohen] for details). If $\Sigma$ is a stable avenue surface, the $\operatorname{Map}(\Sigma)$ is not CB, and so $N=0$ implies that $S$ has finite positive genus.
@@
Since every stable avenue surface is a translatable surface with tame end space, the following result allows us to use the translatable curve graph to understand the large-scale geometry of their mapping class groups.
@@ 3.2. Lassoing a puncture
Given $\alpha$ and $\ell$ as above, the lasso curve defined by $\alpha$ and $\ell$, denoted by $\alpha(\ell)$, is the simple closed curve obtained by tracing along $\alpha$ and $\ell^{\prime}$. In this case, the arc $\ell$ is called a lasso.
To give a more precise definition, consider a small tubular neighborhood of $\alpha\cup\ell^{\prime}$. This subsurface has two boundary components, one of which is isotopic to $\alpha$. Then $\alpha(\ell)$ is equal to the other boundary component.
@@ 3.4. Flux
Let $\alpha$ and $\beta$ be separating curves. The flux of $\alpha$ and $\beta$, denoted by $F(\alpha,\beta)$, is defined to be $F(\alpha,\beta;\gamma)$ for some (any) choice of separating curve $\gamma$ in $\alpha_{+}\cap\beta_{+}$ or $\alpha_{-}\cap\beta_{-}$.
@@
We collect some basic facts about this notion. Lemma 3.11 is used throughout the proof without explicit mention.
@@ 3.5. Main theorem (special case)
For any separating curve $\eta$, there is a mapping class in $\operatorname{Map}(\Sigma)$ which sends $\beta$ to $\eta$ [Schaffer-Cohen, Lemma 3.6]. Therefore, since $\operatorname{Map}(\Sigma)$ acts isometrically on $\Gamma$, we may assume without loss of generality that $\beta$ is a boundary curve of some copy of $S$ in $\Sigma=S^{\natural\mathbb{Z}}$. By Lemma 3.20 (2), $\beta$ is either flux-right or flux-left of $o$. Assume without loss of generality that $\beta$ is flux-right of $o$. Recall that $\Sigma$ admits a translation $h$ such that $h(\beta)$ lies on the right side of $\beta$. Then Lemma 3.20 (6) implies $F(o,h(\beta))>F(o,\beta)$. Moreover, $h(\beta)$ is flux-right of $\beta$ by Lemma 3.20 (1), and so Lemma 3.20 (4) implies that $h(\beta)$ is flux-right of $o$. It follows from induction that $F(o,h^{k}(\beta))\geq F(o,\beta)+k$ for all $k\in\mathbb{N}$. Since $h$ is a translation, we may fix a large enough $k$ so that $h^{k}(\beta)$ lies on the right side of $\alpha$. Put $\beta^{\prime}=h^{k}(\beta)$. Then $(\beta,h(\beta),\dots,h^{k}(\beta))$ is a path from $\beta$ to $\beta^{\prime}$. If this path has length at most $R$, then it must be disjoint from $B$ because $d(\beta,o)=3R$. Otherwise if the length is greater than $R$, then the first $R+1$ vertices on the path lie outside of $B$ for the same reason as before, and the remaining vertices also lie outside of $B$ because for all $R<i\leq k$,
@@
Since $\beta^{\prime}$ lies on the right side of $\alpha$, Lemma 3.18 gives a straight path $(\gamma_{0}=\alpha,\gamma_{1},\dots,\gamma_{n}=\beta^{\prime})$ from $\alpha$ to $\beta^{\prime}$. We now describe a procedure to be iterated several times. Fix a carrier $K$ of $o$ and the $\gamma_{i}$. Let $\ell$ be an arc from $K_{-}$ to a puncture $p\in K_{+}$ such that $\ell$ intersects each $\gamma_{i}$ exactly once, and construct the lasso path $\gamma_{0}(\ell),\dots,\gamma_{n}(\ell)$. As lasso curves, each $\gamma_{i}(\ell_{i})$ is adjacent to $\gamma_{i}$ in $\Gamma$. In particular, $\alpha$ and $\alpha(\ell)$ are connected by an edge, and so are $\beta^{\prime}$ and $\beta^{\prime}(\ell)$. Since $K$ is a carrier containing $o,\alpha,\beta$, Lemma 3.6 implies
@@ 4. Stable avenue surfaces
Now we consider the general setting when $\Sigma$ is a stable avenue surface. Again, we denote by $\Gamma$ the translatable curve graph $\mathcal{T}\mathcal{C}(\Sigma)$ of $\Sigma$. First we need to establish some notation. If a simple closed curve $\eta$ that does not separate $e_{+}$ and $e_{-}$, but $\Sigma\setminus\eta$ does have two components, then one of the components must have neither $e_{+}$ nor $e_{-}$ as an end. The union of this component with $\eta$ is a subsurface of $\Sigma$ with boundary $\eta$, and it is denoted by $C_{\eta}$. If a connected surface $T$ has at least one boundary circle, then denote by $\widehat{T}$ the surface obtained from $T$ by capping a disk onto a boundary circle.
@@ 4.1. General lassoing and flux
Let $\alpha$ be a separating curve, and let $V$ be a clopen subset of ends contained either in ${\mathrm{End}}(\alpha_{+})$ or ${\mathrm{End}}(\alpha_{-})$. By the proof of [Schaffer-Cohen, Lemma 4.6], there exists a simple closed curve $\eta$ disjoint from $\alpha$ which bounds $V$. In particular, $\eta$ does not separate $e_{+}$ and $e_{-}$, and ${\mathrm{End}}(C_{\eta})=V$. Moreover, if $S$ has finite positive genus, then for all $g\geq 0$, $\eta$ may be chosen so that $C_{\eta}$ has genus $g$. Let $\lambda$ be an arc which has one endpoint on $\eta$, is disjoint from $\eta$ otherwise, and intersects $\alpha$. Denote by $\lambda^{\prime}$ the subarc of $\lambda$ obtained by traveling along $\lambda$ from its endpoint on $\eta$ until the first intersection point with $\alpha$. In other words, $\lambda^{\prime}$ is the component of $\lambda\setminus\alpha$ containing the endpoint on $\eta$.
@@
Then put $F_{0}(\alpha,\beta)=p_{0}(\alpha,\beta)+g_{0}(\alpha,\beta)$. Now remove the assumption that $\alpha$ and $\beta$ are disjoint. For separating curves $\gamma$ in $\alpha_{+}\cap\beta_{+}$ or $\alpha_{-}\cap\beta_{-}$, define $p(\alpha,\beta;\gamma)=\left|{p_{0}(\alpha,\gamma)-p_{0}(\beta,\gamma)}\right|$ and $g(\alpha,\beta;\gamma)=\left|{g_{0}(\alpha,\gamma)-g_{0}(\beta,\gamma)}\right|$. The statements and proofs of Lemma 3.11 and Lemma 3.12 all hold for $p_{0}$ and $g_{0}$. So define $p(\alpha,\beta)=p(\alpha,\beta;\gamma)$ and $g(\alpha,\beta)=g(\alpha,\beta;\gamma)$ for some (any) choice of separating curve $\gamma$ in $\alpha_{+}\cap\beta_{+}$ or $\alpha_{-}\cap\beta_{-}$. The flux of separating curves $\alpha$ and $\beta$ is then defined by
@@
Then the same arguments in the previous section may be used to show that $F$ is a pseudometric on $\Gamma$ which bounds $d$ from below. Also, it is still true in this setting that straight paths are geodesic. On the other hand, Lemma 3.18 does not immediately extend to the more general setting. Indeed, even if $\alpha$ and $\beta$ are disjoint, then it is possible that $d(\alpha,\beta)>F_{0}(\alpha,\beta)$. This is due to the subtle definition of $\Gamma$. For example, suppose $\Sigma$ has discrete-type ends, but $\alpha,\beta]$ has genus 1 and no ends. Then $F_{0}(\alpha,\beta)=1$, but $d(\alpha,\beta)\neq 1$ because $\alpha,\beta]$ is not homeomorphic to $T_{n+1}$. However, next we give a sufficient condition on $\alpha,\beta]$ to guarantee that $d(\alpha,\beta)=F(\alpha,\beta)$.

### v2

@@ 2.6. Avenue surfaces and translatable surfaces
Definition 2.7.
An avenue surface is a surface $\Sigma$ which does not contain any non-displaceable subsurface of finite type, whose end space is tame, and whose mapping class group $\operatorname{Map}(\Sigma)$ is CB-generated but not CB.
The condition of tameness only requires that certain ends of $\Sigma$ admit a stable neighborhood. Stable surfaces, in which every end admits a stable neighborhood, automatically have tame end space. Since we will eventually only consider stable surfaces, we refer the interested reader to [MannRafi] for a precise definition of tame and omit writing it here.
Definition 2.8.
Given a surface $\Sigma$, an end $e$ of $\Sigma$, and a sequence $\alpha_{n}$ of simple closed curves on $\Sigma$, we write $\lim_{n\to\infty}\alpha_{n}=e$ if for each neighborhood $V$ of $e$, all but finitely many $\alpha_{n}$ are contained in $V$ (after isotopy). A homeomorphism $h$ of $\Sigma$ is a translation if there are two distinct ends $e_{+}$ and $e_{-}$ of $\Sigma$ such that for any simple closed curve $\alpha$ on $\Sigma$, $\lim_{n\to\infty}h^{n}(\alpha)=e_{+}$ and $\lim_{n\to\infty}h^{-n}(\alpha)=e_{-}$. A surface $\Sigma$ is translatable if it admits a translation.
By [Schaffer-Cohen, Theorem 5.3], if a surface $\Sigma$ has tame end space, and $\operatorname{Map}(\Sigma)$ is CB generated and not CB, then $\Sigma$ is an avenue surface if and only if it is translatable. In particular, every stable avenue surface is translatable.
@@
Next we recall a characterization of translatable surfaces from [Schaffer-Cohen]. Let $S$ be any surface with exactly two boundary components, and denote by $S^{\natural\mathbb{Z}}$ the surface obtained by arranging countably many copies of $S$ like $\mathbb{Z}$ and then gluing consecutive copies of $S$ together along their boundary components in the natural way. By construction, this surface is translatable. Conversely, every translatable surface admits such a decomposition.
@@ 2.7. The translatable curve graph
We now recall the definition of the translatable curve graph from [Schaffer-Cohen]. Let $\Sigma=S^{\natural\mathbb{Z}}$ be a translatable surface and assume that $\Sigma$ is tame. By [Schaffer-Cohen, Lemma 4.4] and [MannRafi, Proposition 4.7], ${\mathrm{End}}(S)$ has a finite, positive number of equivalence classes of maximal ends, and each such equivalence class intersects ${\mathrm{End}}(S)$ at either a finite set or a Cantor set. So we may fix, once and for all, a set $\mathcal{R}=\left\{{f_{1},\dots,f_{N},c_{1},\dots,c_{M}}\right\}$ of representatives of the equivalence classes of maximal ends of $S$ such that each $E(f_{i})\cap{\mathrm{End}}(S)$ is finite and each $E(c_{i})\cap{\mathrm{End}}(S)$ is a Cantor set. Then in the whole end space ${\mathrm{End}}(\Sigma)$, since $\Sigma=S^{\natural\mathbb{Z}}$, each $E(f_{i})$ is homeomorphic to $\mathbb{Z}$ and each $E(c_{i})$ is homeomorphic to a Cantor set with two points removed. We say an end of $\Sigma$ is of discrete type if it is equivalent to some $f_{i}$. Observe that $\Sigma$ has at least one end of discrete type if and only if $N>0$.
@@
Pick mutually disjoint stable neighborhoods $V_{f_{i}}$ and $V_{c_{i}}$ of the elements of $\left\{{f_{1},\dots,f_{N},c_{1},\dots,c_{M}}\right\}$. The definition of stable implies that each $V_{f_{i}}$ contains exactly one discrete-type end. For each $i=1,\dots,N$, let $T_{i}$ be a subsurface of $S$ with two boundary circles and end space homeomorphic to $V_{f_{i}}\sqcup\bigsqcup_{j=1}^{M}V_{c_{j}}$. If $f_{i}$ or any of the $c_{j}$ is accumulated by genus, then $T_{i}$ necessarily has infinite genus; otherwise, we require $T_{i}$ to have genus 0. If $S$ has finite positive genus, then let $T_{N+1}$ be a subsurface with two boundary circles, genus 1, and end space homeomorphic to $\bigsqcup_{j=1}^{M}V_{c_{j}}$. Put $\mathcal{S}=\left\{{T_{1},\dots,T_{N},T_{N+1}}\right\}$ (including $T_{N+1}$ only if it is defined, of course). The translatable curve graph $\mathcal{T}\mathcal{C}(\Sigma)$ of $\Sigma$ is the graph whose vertices are the isotopy classes of separating curves on $\Sigma$, with an edge between two curves $\alpha$ and $\beta$ if they have disjoint representatives and $\alpha,\beta]$ is homeomorphic to a subsurface in $\mathcal{S}$. By [Schaffer-Cohen, Lemma 4.7], $\mathcal{T}\mathcal{C}(\Sigma)$ is connected, and we denote its path metric by $d$. Observe that $\operatorname{Map}(\Sigma)$ naturally acts by isometries on $\mathcal{T}\mathcal{C}(\Sigma)$. Furthermore, this action is transitive [Schaffer-Cohen, Lemma 3.6].
@@
There is an edge case: if $N=0$ and the genus of $S$ is either 0 or infinite, then the above construction gives $\mathcal{S}=\emptyset$. This is not desirable, so in this case we put $\mathcal{S}=\{S\}$. In some sense this choice is consistent with the construction above because $S$ is indeed a subsurface with two boundary circles and end space homeomorphic to $\bigsqcup V_{c_{i}}$. More importantly, in this case $\mathcal{T}\mathcal{C}(\Sigma)$ has diameter 2 and $\operatorname{Map}(\Sigma)$ is CB, which means that both are quasi-isometric to a point (see [Schaffer-Cohen] for details). If $\Sigma$ is a stable avenue surface, then $\operatorname{Map}(\Sigma)$ is not CB by assumption, and so $N=0$ implies that $S$ has finite positive genus.
@@
Since every stable avenue surface $\Sigma$ is a translatable surface with tame end space, the following result allows us to use the translatable curve graph to understand the large-scale geometry of $\operatorname{Map}(\Sigma)$.
@@ 3.2. Lassoing a puncture
Given $\alpha$ and $\ell$ as above, the lasso curve defined by $\alpha$ and $\ell$, denoted by $\alpha(\ell)$, is the simple closed curve obtained by tracing along $\alpha$ and $\ell^{\prime}$. In this case, the arc $\ell$ is called a lasso. See Figure 3.1.
Figure 3.1. Lassoing a puncture: $\alpha$ is the black curve in the middle, $\ell$ is the blue arc, $\alpha(\ell)$ is the new red curve.
To give a more precise definition, consider a small tubular neighborhood of $\alpha\cup\ell^{\prime}$. This subsurface has two boundary components, one of which is isotopic to $\alpha$. Then $\alpha(\ell)$ is equal to the other boundary component. The terminology is intended to evoke the image of the curve $\alpha$ using the arc $\ell$ as a lasso to wrangle the puncture $p$.
@@ 3.3. Lasso paths
See Figure 3.2 for an example of a straight path and a lasso path.
@@ 3.4. Flux
Let $\alpha$ and $\beta$ be separating curves. The flux between $\alpha$ and $\beta$, denoted by $F(\alpha,\beta)$, is defined to be $F(\alpha,\beta;\gamma)$ for some (any) choice of separating curve $\gamma$ in $\alpha_{+}\cap\beta_{+}$ or $\alpha_{-}\cap\beta_{-}$.
@@
We collect some basic facts about this notion. Lemma 3.11 is used throughout the proof of the next lemma without explicit mention.
@@ 3.5. Main theorem (special case)
Since $\operatorname{Map}(\Sigma)$ acts transitively on $\Gamma$, we may assume without loss of generality that $\beta$ is a boundary curve of some copy of $S$ in $\Sigma=S^{\natural\mathbb{Z}}$. By Lemma 3.20 (2), $\beta$ is either flux-right or flux-left of $o$. Assume without loss of generality that $\beta$ is flux-right of $o$. Let $h$ be the translation on $\Sigma$ such that $h(\beta)$ lies on the right side of $\beta$ and $\beta,h(\beta)]$ is homeomorphic to $S$. Then Lemma 3.20 (6) implies $F(o,h(\beta))>F(o,\beta)$. Moreover, $h(\beta)$ is flux-right of $\beta$ by Lemma 3.20 (1), and so Lemma 3.20 (4) implies that $h(\beta)$ is flux-right of $o$. It follows from induction that $F(o,h^{i}(\beta))\geq F(o,\beta)+i$ for all $i\in\mathbb{N}$. Since $h$ is a translation, we may fix a large enough $k$ so that $h^{k}(\beta)$ lies on the right side of $\alpha$. Put $\beta^{\prime}=h^{k}(\beta)$. Then $(\beta,h(\beta),\dots,h^{k}(\beta))$ is a path from $\beta$ to $\beta^{\prime}$. If this path has length at most $R$, then it must be disjoint from $B$ because $d(\beta,o)=3R$. Otherwise if the length is greater than $R$, then the first $R+1$ vertices on the path lie outside of $B$ for the same reason as before, and the remaining vertices also lie outside of $B$ because for all $R<i\leq k$,
@@
Since $\beta^{\prime}$ lies on the right side of $\alpha$, Lemma 3.18 gives a straight path $(\gamma_{0}=\alpha,\gamma_{1},\dots,\gamma_{n}=\beta^{\prime})$ from $\alpha$ to $\beta^{\prime}$. We now describe a procedure to be iterated several times. Fix a carrier $K$ of $o$ and the $\gamma_{i}$. Let $\ell$ be an arc from $K_{-}$ to a puncture $p\in K_{+}$ such that $\ell$ intersects each $\gamma_{i}$ exactly once, and construct the lasso path $\gamma_{0}(\ell),\dots,\gamma_{n}(\ell)$. See Figure 3.2 for an example. As lasso curves, each $\gamma_{i}(\ell_{i})$ is adjacent to $\gamma_{i}$ in $\Gamma$. In particular, $\alpha$ and $\alpha(\ell)$ are connected by an edge, and so are $\beta^{\prime}$ and $\beta^{\prime}(\ell)$. Since $K$ is a carrier containing $o,\alpha,\beta^{\prime}$, Lemma 3.6 implies
@@
Figure 3.2. The lasso path constructed from a geodesic between $\alpha$ and $\beta^{\prime}$ and an arc $\ell$ ending at a puncture outside a carrier $K$.
@@ 4. Stable avenue surfaces
Now we consider the general setting when $\Sigma$ is a stable avenue surface. Again, we denote by $\Gamma$ the translatable curve graph $\mathcal{T}\mathcal{C}(\Sigma)$ of $\Sigma$. First we need to establish some notation. If a simple closed curve $\eta$ in $\Sigma$ does not separate $e_{+}$ and $e_{-}$, but $\Sigma\setminus\eta$ does have two components, then one of the components must have neither $e_{+}$ nor $e_{-}$ as an end. The union of this component with $\eta$ is a subsurface of $\Sigma$ with boundary $\eta$, and it is denoted by $C_{\eta}$. If a connected surface $T$ has at least one boundary circle, then denote by $\widehat{T}$ the surface obtained from $T$ by capping a disk onto a boundary circle.
@@ 4.1. General lassoing and flux
Let $\alpha$ be a separating curve in $\Sigma=S^{\natural\mathbb{Z}}$, and let $V$ be a clopen subset of ends contained either in ${\mathrm{End}}(\alpha_{+})$ or ${\mathrm{End}}(\alpha_{-})$. By the proof of [Schaffer-Cohen, Lemma 4.6], there exists a simple closed curve $\eta$ disjoint from $\alpha$ which bounds $V$. In particular, $\eta$ does not separate $e_{+}$ and $e_{-}$, and ${\mathrm{End}}(C_{\eta})=V$. Moreover, if $S$ has finite positive genus, then for all $g\geq 0$, $\eta$ may be chosen so that $C_{\eta}$ has genus $g$. Let $\lambda$ be an arc which has one endpoint on $\eta$, is disjoint from $\eta$ otherwise, and intersects $\alpha$. Denote by $\lambda^{\prime}$ the subarc of $\lambda$ obtained by traveling along $\lambda$ from its endpoint on $\eta$ until the first intersection point with $\alpha$. In other words, $\lambda^{\prime}$ is the component of $\lambda\setminus\alpha$ containing the endpoint on $\eta$.
@@
Then put $F_{0}(\alpha,\beta)=p_{0}(\alpha,\beta)+g_{0}(\alpha,\beta)$. Now remove the assumption that $\alpha$ and $\beta$ are disjoint. For separating curves $\gamma$ in $\alpha_{+}\cap\beta_{+}$ or $\alpha_{-}\cap\beta_{-}$, define $p(\alpha,\beta;\gamma)=\left|{p_{0}(\alpha,\gamma)-p_{0}(\beta,\gamma)}\right|$ and $g(\alpha,\beta;\gamma)=\left|{g_{0}(\alpha,\gamma)-g_{0}(\beta,\gamma)}\right|$. The statements and proofs of Lemma 3.11 and Lemma 3.12 all hold for $p_{0}$ and $g_{0}$. So define $p(\alpha,\beta)=p(\alpha,\beta;\gamma)$ and $g(\alpha,\beta)=g(\alpha,\beta;\gamma)$ for some (any) choice of separating curve $\gamma$ in $\alpha_{+}\cap\beta_{+}$ or $\alpha_{-}\cap\beta_{-}$. The flux between $\alpha$ and $\beta$, denoted by $F(\alpha,\beta)$, is then defined by
@@
Then the same arguments in the previous section may be used to show that $F$ is a pseudometric on $\Gamma$ which bounds $d$ from below. Also, it is still true in this setting that straight paths are geodesic. On the other hand, Lemma 3.18 does not immediately extend to the more general setting. Indeed, even if $\alpha$ and $\beta$ are disjoint, then it is possible that $d(\alpha,\beta)>F_{0}(\alpha,\beta)$. This is due to the subtle definition of $\Gamma$. For example, suppose $\Sigma=S^{\natural\mathbb{Z}}$ and $S$ has a maximal end which is not of discrete type. If $\alpha,\beta]$ has genus 1 and no ends, then $F_{0}(\alpha,\beta)=1$, but $d(\alpha,\beta)\neq 1$ because $\alpha,\beta]$ is not homeomorphic to $T_{N+1}$. However, we give a sufficient condition on $\alpha,\beta]$ to guarantee that $d(\alpha,\beta)=F(\alpha,\beta)$.